Hindelang Inc. is considering a project that has the following cash flow and WACC data. What is the project's MIRR? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box. WACC: 11% Year: 0 1 2 3 4 Cash flows: -$875 $350 $375 $400 $425
Answer #1

Let Cash flow for period n be CFn

WACC = r = 12.35%

Given,

CF0 = -875
CF1 = 350
CF2 = 375
CF3 = 400
CF4 = 425

MIRR = ( FVc / PVfc )1/n -1

where, FVc is the  future value of positive cash flows

and PVfc is the present value of negative cash flows

PVfc = 875

FVc = 350*1.12353 + 375*1.12352 + 400*1.1235 + 425 = 1844.09

=> MIRR = (1844.09/875)1/4 - 1 = 0.2049 = 20.49%